OpenStudy (anonymous):

simplify the equation 3x / 21(x+y)

jimthompson5910 (jim_thompson5910):

fill in the blank 21 = 3 times ______

OpenStudy (anonymous):

7 i know but the only problem is the x and y

jimthompson5910 (jim_thompson5910):

so that means \[\Large \frac{3x}{21(x+y)}\] \[\Large \frac{3*x}{3*7(x+y)}\]

jimthompson5910 (jim_thompson5910):

what term do you see cancelling?

OpenStudy (anonymous):

the 3 thats is so would it be 1/7(x+y)

jimthompson5910 (jim_thompson5910):

good \[\Large \frac{3x}{21(x+y)}\] \[\Large \frac{3*x}{3*7(x+y)}\] \[\Large \frac{\cancel{3}*x}{\cancel{3}*7(x+y)}\] \[\Large \frac{x}{7(x+y)}\]

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

You can stop there or you can optionally distribute to go from \[\Large \frac{x}{7(x+y)}\] to \[\Large \frac{x}{7x+7y}\]

jimthompson5910 (jim_thompson5910):

yw